OpenStudy (anonymous):

I have another log question :o Log3(x^2)=2log3(4) - 4log3(5)

OpenStudy (irishboy123):

latexed \(log_3(x^2)=2log_3(4) - 4log_3(5)\) is that right?

OpenStudy (anonymous):

Yep, that's right!

OpenStudy (welshfella):

by the laws of logs log3 x^2 = log3 4^2 - log3 4^5 log3 x^2 = log3 (4^2 / 4^5) can you continue?

OpenStudy (anonymous):

How did you get 4^5? Isn't it 5^4?

OpenStudy (welshfella):

** ooops sorry that is 5^4 not 4^5

OpenStudy (anonymous):

I got to where you are, and I understand it, but now I'm unsure of how to continue from there :o

OpenStudy (welshfella):

this is easy now..

OpenStudy (anonymous):

:/ Still not sure of how to proceed, sorry.

OpenStudy (welshfella):

if a^x = a^y what is the relation between x and y?

OpenStudy (anonymous):

They're equal to each other, right?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

x^2 = 16/625?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

so the final answer is x=4/25, right? Just want to make sure.

OpenStudy (welshfella):

right - what about -4/25?

OpenStudy (welshfella):

catch question..

OpenStudy (anonymous):

Since it's log, you can't have a negative answer..

OpenStudy (anonymous):

...right?

OpenStudy (irishboy123):

why not put \(-\frac{4}{25}\) into the original equation....? see if it works you cruel man, @welshfella !!

OpenStudy (welshfella):

i know!!